OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = , x = − x = −1, x = 9 x = −1, x = −9 No solution

OpenStudy (jdoe0001):

\(\bf -4|x + 5| = -16\implies |x + 5| =\cfrac{-16}{-4}\\ \quad \\ |x + 5| =\cfrac{-16}{-4}\implies \begin{cases} +(x + 5) =\cfrac{-16}{-4}\\ \quad \\ \bf -(x + 5) =\cfrac{-16}{-4} \end{cases}\)

OpenStudy (jdoe0001):

so you'd get 2 values for "x", one for each scenario

OpenStudy (anonymous):

x=-1